OpenStudy (anonymous):

Compare the weight of a person standing on the equator to the centripetal force experienced by that person due to the Earth's rotation. The results will be used to dispel that people on the equator feel lighter due to the rotation of the Earth. - You may use any mass for a person in your calculations. You must clearly show how you calculated the magnitudes of both forces and how they compare to each other. Explain whether your answers depend on the mass of the person or would anyone feel the same way. You should state whether the centripetal force makes a significant impact on the weight.

OpenStudy (anonymous):

Here is what I got: The centripetal acceleration at the equator = 4*pi^2*r/T^2 Where T = time period. For Earth, this is 24hours = 86400 seconds. R = radius of Earth. This is 6400 km = 6400000 meters Therefore, 4*pi^2*6400000/86400^2 = 0.03 m/s^2 Since the acceleration due to gravity is approx. 9.81m/s^2, then you would weigh only about 0.3% less at the equator than at either the North or South poles. Is that right? I feel like I am missing something. Any help would be much appreciated! Thanks so much!

OpenStudy (anonymous):

Must be too challenging. :p

OpenStudy (anonymous):

Any help? :(

OpenStudy (anonymous):

Well.. That was fun.

OpenStudy (anonymous):

Right idea. where did pi^2 come from? should just be pi? You want to compare v^2/r acceleration with g=9.8m/s^2. v = 2 pi r / T.

OpenStudy (anonymous):

I figured this out already. I appreciate your effort to help nonetheless1